kward64:

what is 124 divided by 1,860

BeeBoy:

0.0666

kward64:

@beeboy wrote:
0.0666
not in a decimal

BeeBoy:

0

Extrinix:

\(\dfrac{124}{1860}\) would convert to roughly 0.1, but as a fraction would be \(\dfrac{1}{15}\) Sorry for the direct answer, there's no real explaining this one x.x

kward64:

not as a fraction

Extrinix:

The only answers are fractions or decimals

kward64:

nvm i figured it out

Extrinix:

Unless you mean \(\dfrac{1860}{124}\), which means you made your wording wrong \(\dfrac{1860}{124}\) would be \(15\) if you meant it that way ?

BeeBoy:

you could have just looked it up -

Extrinix:

@beeboy wrote:
you could have just looked it up -
^^^

kward64:

yeah thats how i meant it

kward64:

just nvm

Extrinix:

Ah okay, word it correctly next time please

BeeBoy:

i said 0 earlier lol '

TETSXPREME:

What is 124*__=

kward64:

@beeboy wrote:
i said 0 earlier lol '
but its not zero

BeeBoy:

I know this

kward64:

@tetsxpreme wrote:
What is 124*__=
i already figured it out

kward64:

@kward64 wrote:
@tetsxpreme wrote:
What is 124*__=
i already figured it out
thanks tho